Wires A, B, and C meet at a junction. The current in wire A is 3.8 mA into the junction, and the current in wire B is 5.7 mA out of the junction.

What is the current in wire C, and in what direction, into or out of the junction, is this current?

Express your answer with the appropriate units. Enter positive value if the current is into the junction and negative value if the current is out of the junction.
Answer #1

To obtain the answer of this we will apply Kirchoff's Current Law, which states that at a junction, the net current entering the junction equals the net current exiting the junction.

So, the sum of the currents from each of the wires should equate to zero. Considering current into the junction as positive and out of the junction as negative,

Let IA = Current in wire A = 3.8 mA (As current is entering the junction)

IB = Current in wire B = = - 5.7 mA ( As current from this wire is leaving the junction)

IC = Current in wire C

Now IA + IB + IC = 0

So, IC = - IA -IB  = -3.8 -(-5.7) mA = 1.9 mA

Hence Current flowing through wire C is 1.9 mA and positive sign indicates, current is flowing into the junction.